The average velocity of an ideal gas molecule at 27°C is 0.3 m/s. The average velocity at 927°C will be
a. 0.6 m/s
b. 0.3 m/s
c. 0.9 m/s
d. 3.0 m/s

Grade:11

1 Answers

Navjyot Kalra
9 years ago
Sol. Expression for average velocity is uavg = √(8RT/πM) For the same gas but at different temperature (u_avg 〖(T〗_1) )/(u_avg (T_2 ) ) = √(T_1/T_2 ) = √(300/1200) = 1/2 ⇒ uavg (927°C) = 2 × uavg (27°C) = 0.6 ms-1